OpenStudy (anonymous):

the second of two numbers is 1 more than twice the first. their sum is 25. find the numbers

OpenStudy (shadowfiend):

So what are the two equations we can get from this?

OpenStudy (anonymous):

y=x+2+1 ?? x+y+25

OpenStudy (anonymous):

i mean X*2+1

OpenStudy (shadowfiend):

Correct (and I assume you meant x + y = 25).

OpenStudy (anonymous):

yes. lol

OpenStudy (shadowfiend):

Ok, so how can you solve this system?

OpenStudy (anonymous):

x+x*2+1=25

OpenStudy (anonymous):

5x=25

OpenStudy (shadowfiend):

Hm. Where did you get 5x from?

OpenStudy (anonymous):

this is where i am having a prob. I get confused on what to add or multiply first

OpenStudy (shadowfiend):

The easiest way is to group like terms (e.g, terms with x) and then go left to right. For example, in this case they're already grouped: x + 2x + 1 = 25 So we ask ourselves, what is x + 2x?

OpenStudy (anonymous):

3x+1=25 ??

OpenStudy (shadowfiend):

Right. So then what do you do to bring the 1 to the right?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

x=8

OpenStudy (shadowfiend):

Perfect :)

OpenStudy (anonymous):

y=17

OpenStudy (anonymous):

how do I know if i am solving for x or y?

OpenStudy (shadowfiend):

Right on.

OpenStudy (shadowfiend):

You have to solve for both. For the first step, you just have to pick one.

OpenStudy (anonymous):

I have a hard time with setting it up. sometimes I do X and sometimes it's y

OpenStudy (shadowfiend):

As long as you make sure to keep referring to the same variable as x and y throughout, your result should make sense at the end. You can check it by going out to your very first equation and plugging in your answers and making sure that the numbers on both sides are actually equal.

OpenStudy (anonymous):

thank you

OpenStudy (shadowfiend):

No problem.